# references.bib

### v3

@@ I Introduction
Communication for synchronizing embedding lookups and gradients across all GPUs during each minibatch iteration significantly contributes to the overall training time. For example, Figure 1 shows that all-to-all communication accounts for more than 60% of the total training time for DLRM on an 8-node, 32 A100 GPUs cluster (connected through a Slingshot 10 interconnect [DBLP:conf/sc/SensiGMRH20]). Consequently, various studies have been conducted to address these communication challenges.
@@ II-A Deep Learning Recommendation Model (DLRM)
DLRM Architecture Generally, DLRM comprises three components: Embedding, MLP, and Interaction Module. The architecture is shown in Figure 2. Embedding tables in DLRM process categorical data, mapping each categorical feature into a vector representation in an embedding vector space. There are two MLPs in DLRM: the Bottom MLP and the Top MLP. The Bottom MLP transforms continuous features to match the length of embedding vectors. The Top MLP applies the concatenated data as input and calculates the Click-Through Rate (CTR) as output.
DLRM Training Pipeline and Bottlenecks Training DLRM involves both model and data parallelism to manage its diverse computational needs efficiently. Model parallelism is crucial for handling the large Embedding Tables (EMBs) distributed across different devices due to their size, while data parallelism is applied to the Multilayer Perceptrons (MLPs), which, despite requiring full access every epoch, consume a relatively small amount of memory. This dual approach allows for the division of a global batch of embedding vectors into smaller local batches, facilitating interaction operations and subsequent processing by the Top MLP through an all-to-all communication pattern. This setup is particularly effective for implementing various second-order interaction methods, including dot products between pairs of embedding vectors and dense features, which are then concatenated with dense features for input to the Top MLP, ultimately enabling classification. During backward propagation, a symmetrical all-to-all communication redistributes gradients back to their respective devices for updating the Embedding tables and Bottom MLP, reflecting the forward phase’s operations.
@@ II-B Floating-point Data Compression
A new generation of lossy compression techniques for scientific data, particularly floating-point data, has emerged, exemplified by SZ [sz16, sz17, sz3liang], ZFP [zfp], MGARD [mgard], and TTHRESH [BLP:19]. Unlike their counterparts for media, these scientific data compressors offer strict error-controlling schemes, enabling users to manage accuracy loss both in the reconstructed data and during post-analysis.
With the proliferation of GPU-based HPC systems and applications, compressors like SZ, ZFP, and MGARD have introduced GPU-optimized versions (i.e., cuSZ [cusz2020], cuZFP [zfp], and MGARD-GPU [chen2021accelerating]), delivering significantly enhanced throughput compared to their CPU-based implementations. ZFP, a transform-based compressor, allows users to set a desired bitrate, while SZ, a prediction-based compressor, and MGARD, a multigrid-based compressor, enable the specification of a maximum tolerable error. ZFP in fixed-rate mode tends to offer consistently higher throughput, whereas SZ and MGARD in error-bounded mode achieve superior compression ratios.
@@ II-C Problem Statement
Dominance of Communication Overhead. As illustrated in Figure 1, all-to-all communication accounts for over 60% of DLRM’s total training time, establishing communication as the bottleneck rather than computation. This bottleneck is exacerbated when training DLRM with datasets of various sizes. For instance, the Criteo Kaggle dataset, with sparse feature lengths of 32, generates more than 121 GB of lookup data per epoch. This figure escalates dramatically with larger datasets, such as the Criteo Terabyte dataset, which can accumulate up to 26 TB of lookup data per epoch. The scale increases further with industry-level recommendation models, often exceeding multiple terabytes, necessitating larger volumes of training data and larger-scale distributed systems for parameter storage. This significantly increases communication data across devices, highlighting the urgent need to reduce communication data volume in DLRM training.
However, as aforementioned, there are several challenges to address this issue due to the limitations of directly applying current error-bounded lossy compression techniques. To effectively tackle these hurdles, this paper focuses on the following key research questions: ❶ Error Bound Configuration: It’s essential to determine the optimal error bounds for various embedding tables across different training iterations to maintain training accuracy while enhancing compression ratios. ❷ DLRM-specific Lossy Compression Algorithm: There’s a critical need to devise a compression algorithm uniquely suited for DLRM’s embedding tables, aiming for elevated compression ratios without substantially compromising data integrity. ❸ GPU Compression Performance Optimization: Optimizing the GPU execution of our specialized compression and ensuring its smooth incorporation within DLRM training is crucial for improving overall training performance.
@@ III-A Overview of DLRM Training Pipeline with Compression
Offline Analysis. The purpose of this step is to obtain an optimized configuration by sampling and analyzing some iterations from the original training process. There are two tasks involved in offline analysis: Compressor Selection and Embedding Table Classification. In the Compressor Selection task, we evaluate various compressors using sampled data to select the best one for the current system. In the Embedding Table Classification task, we analyze the characteristics of embedding tables using sampled data and classify them according to different error bound adjustment strategies.
@@ III-B Observation and In-Depth Analysis of Embedding Vector
TABLE I: Characteristics of thre representative EMB tables from Criteo Kaggle dataset.
@@
Violently Vector Homogenization $\checkmark$ $\times$ $\times$
Gaussian Distribution $\checkmark$ $\checkmark$ $\times$
@@ III-C Adaptive Fine-Grain Error-Bound Adjust Strategy
Specifically, we divide the training period into two phases: an initial phase, characterized by rapid loss reduction, and a later phase, where the loss tends to stabilize. During the initial phase, we gradually decrease the error bound via a predefined decay function (e.g., logarithmic, stepwise) to minimize deviation from the original data, facilitating swift model convergence. In the later phase, we maintain a consistent error bound to ensure the model converges effectively.
@@
Figure 6: EMB Table size of Criteo Kaggle and Criteo Terabytes Dataset
@@
$\eta=\frac{N_{original}-N_{compressed}}{N_{original}}.$ (1)
@@
Overall, we detail our proposed adaptive error bound adjustment strategy, as outlined in Listing 1.
Listing 1: Proposed Error Bound Adjustment Strategy
@@
if c $\in$ L_EMB: EBConf[t] $\leftarrow$ LargeEB
if c $\in$ S_EMB: EBConf[t] $\leftarrow$ SmallEB
@@ III-D Optimized Compression Algorithm for DLRM Data
After identifying the optimal error bound for training, the next step involves fine-tuning the compressor within that error bound. Compressors typically aim to balance achieving a high compression ratio with maintaining high compression speed. The ideal compressor for our purposes should satisfy three criteria: ❶ Operate on the GPU to avoid data transfers between the device and host, since embedding vectors reside on the GPU. ❷ Offer high compression throughput to minimize compression overhead and, consequently, accelerate overall DLRM training. ❸ Achieve a high compression ratio to maximize the benefits of reduced data volume during communication. To meet these requirements, we propose an optimized hybrid error-bounded lossy compression algorithm tailored specifically for DLRM data.
@@
$speedup=\frac{1}{\frac{1}{CR}+B\times(\frac{1}{T_{c}}+\frac{1}{T_{d}})}$ (2)
Overall, in our hybrid compression framework, we use this formula to pinpoint the most efficient compressor that maximizes speed-up for the given system configuration. Although theoretically any compression algorithm could be included in our selection pool, for simplicity and effectiveness, we limit our final design to these two encoders.
Listing 2: Compressor Selection
@@
spd $\leftarrow$ SpeedUpCompute(d, Compressors)
@@ III-E System Implementation and Performance Optimization
Compressor Fine-tuning. First, we refine our vector-based LZ compression algorithm, distinguishing it from the original LZ approach in two significant ways. ❶ Extended window size: Traditional LZ algorithms typically opt for a relatively small window size, like 4KB/8KB. However, given the longer repeated patterns observed in DLRM applications, with lengths ranging from 128-256 bytes (considering an embedding vector length of 32/64 with 32-bit floating point data type), we extend the window size to accommodate these patterns. ❷ Fixed pattern length: To minimize memory access and byte comparison time, we introduce fixed-length pattern matching. In contrast to the standard LZ process, which advances the pointer to the next byte in the absence of a match, our knowledge of the repeated pattern length allows us to leap forward many bytes in search of the next match, thereby avoiding fruitless comparisons and pointer movements. If the initial bytes of two embedding vectors differ, further comparison is unnecessary. This optimization also forces the compressor to match longer patterns rather than shorter patterns. Compared to the current state-of-the-art (SOTA) GPU lossless compressor nvCOMP-LZ4, our approach achieves 2.72$\times$ and 4.88$\times$ higher compression ratios on two dataset, respectively.
Buffer Optimization. Next, we implement multi-threading to minimize memory copy overhead and speed up decompression. Typically, compressors output compressed data to a memory chunk and return a pointer, a versatile but overhead-inducing solution. Given the all-to-all collective communication in DLRM training, where each data chunk must be compressed separately for transmission to each rank, this method introduces unnecessary memory copying since compressed data may not be stored contiguously. Besides, launching kernel multiple times introduces overhead of a lot of kernel launching. To address these, we optimize our compressor that not only introduces one kernel launching but also writes directly to the sending buffer. In the compression process, we involve synchronization and Atomic Add to get the writing offset of each chunk. Furthermore, we leverage multi-threading for compression and decompression, partitioning the compressed data into multiple chunks for simultaneous processing. Although GPU compressors typically achieve high resource utilization, some operations cannot fully leverage GPU resources. Thus, executing multiple decompression kernels in parallel is faster than serial decompression.
Figure 7: Proposed compression and decompression optimization.
@@ IV-A Experimental Setup
TAB. ID EB. # Ori. Patterns # Quant. Patterns Batch Size Homo Index
@@
TAB. ID EB. # Ori. Patterns # Quant. Patterns Batch Size Homo Index
@@ IV-B Evaluation of Error Bound Adjustment Strategy
To assess the effects of lossy compression on model accuracy, we evaluate the model prediction accuracy of our training method compared with original DLRM training using both FP32 and FP16 precision, alongside a SOTA low precision approach employing 8-bit quantization. Current standards deem an accuracy loss within 0.02% as acceptable in production models [0.02threshold].
Figure 9(a) displays a comparison of accuracy convergence curves across these methods. Through extensive experimentation, we determine and apply a suitable fixed global error bound (i.e., 0.02) for each model to guarantee convergence. On the two datasets we utilized, the average prediction accuracy losses are 0.0031% and 0.0042%, respectively.
@@
As illustrated in Table II, embedding tables are classified into three categories according to their Homogenization Index scores, with assigned error bounds of 0.01, 0.03, and 0.05, respectively. The accuracy evaluation, detailed in Figure 10(a), reveals that our approach, which applies specific error bounds to different tables rather than a uniform global error bound, maintains the model’s accuracy intact. Additionally, this method achieves a higher compression ratio, up to 1.21$\times$ on Criteo Kaggle dataset, compared to the fixed global error bound strategy.
@@
Figure 10: Accuracy and compression ratio of our method with proposed table-wise EB configuration strategy on different datasets.
@@
Our experiments demonstrated that a step-wise (staircase descent) decay function offers the greatest compression benefits while ensuring model convergence, as shown in Figure 5(b). This finding led us to employ the step-wise function for comparison against the aggressive adjustment approach. Comparative experiments illustrated in Figure 11 reveal how the model training and compression ratios are affected when the error bound is reduced from twice and three times the conservative error bound down to the conservative error bound.
@@
(a) Accuracy
(b) Compression Ratio
Figure 11: Accuracy and compression ratio of our method with two decay methods on different datasets. Decay__\__2x represents the error bound decayed from 2 times base error bound to base error bound; Drop__\__2x represents the error bound dropped from 2 times base error to base error bound at the end of the initial phase.

### v4

# Accelerating Communication in Deep Learning Recommendation Model Training with Dual-Level Adaptive Lossy Compression

@@ I Introduction
Communication for synchronizing embedding lookups and gradients across all GPUs during each mini-batch iteration significantly contributes to the overall training time. For example, Figure 1 shows that all-to-all communication accounts for more than 60% of the total training time for DLRM on an 8-node, 32 A100 GPUs cluster (connected through a Slingshot 10 interconnect [DBLP:conf/sc/SensiGMRH20]). Consequently, various studies have been conducted to address these communication challenges.
@@ II-A Deep Learning Recommendation Model (DLRM)
DLRM Architecture Generally, DLRM comprises three components: Embedding tables, Interaction Module, and MLP. The architecture is shown in Figure 2. Embedding tables in DLRM process categorical data by looking up each categorical feature and mapping it into an embedding vector representation as the output vector. The Interaction Module will apply input vectors from the embedding tables and the Bottom MLP, and perform interactions on them to generate a new output vector. There are two MLPs in DLRM: the Bottom MLP and the Top MLP. The Bottom MLP transforms dense features to match the length of embedding vectors. The Top MLP applies the concatenated data as input and calculates the Click-Through Rate (CTR) as output.
DLRM Training Parallelisms and Bottlenecks Training DLRM involves both model and data parallelism to manage its diverse computational needs efficiently. Model parallelism is crucial for handling the large Embedding Tables (EMBs) distributed across different devices due to their size, while data parallelism is applied to the Multilayer Perceptrons (MLPs), which, despite requiring full access every epoch, consume a relatively small amount of memory. This dual approach allows for the division of a global batch of embedding vectors into smaller local batches, facilitating interaction operations and subsequent processing by the Top MLP through an all-to-all communication pattern. This setup is particularly effective for implementing various second-order interaction methods, including dot products between pairs of embedding vectors and dense features, which are then concatenated with dense features for input to the Top MLP, ultimately enabling classification. During backward propagation, a symmetrical all-to-all communication redistributes gradients back to their respective devices for updating the Embedding tables and Bottom MLP, reflecting the forward phase’s operations.
@@ II-B Floating-point Data Compression
A new generation of lossy compression techniques for scientific data, particularly floating-point data, has emerged, exemplified by SZ [sz16, sz17, sz3liang], ZFP [zfp], and TTHRESH [BLP:19]. Unlike their counterparts for media, these scientific data compressors offer strict error-controlling schemes, enabling users to manage accuracy loss both in the reconstructed data and during post-analysis.
With the proliferation of GPU-based HPC systems and applications, compressors like SZ and ZFP have introduced GPU-optimized versions (i.e., cuSZ [cusz2020] and cuZFP [zfp]), delivering significantly enhanced throughput compared to their CPU-based implementations. ZFP, a transform-based compressor, allows users to set a desired bitrate, while SZ, a prediction-based compressor, enables the specification of a maximum tolerable error. ZFP in fixed-rate mode tends to offer consistently higher throughput, whereas SZ in error-bounded mode achieves superior compression ratios.
@@ II-C Problem Statement
Dominance of Communication Overhead. As illustrated in Figure 1, all-to-all communication accounts for over 60% of DLRM’s total training time, establishing communication as the bottleneck rather than computation. This bottleneck is exacerbated when training DLRM with datasets of various sizes. For instance, the Criteo Kaggle dataset, with sparse feature lengths of 32, generates more than 121 GB of lookup data per epoch. This figure escalates dramatically with larger datasets, such as the Criteo Terabyte dataset, which can accumulate up to terabytes of lookup data per epoch. The scale increases further with industry-level recommendation models, often exceeding multiple terabytes, necessitating larger volumes of training data and distributed systems of larger scales for parameter storage. This significantly increases communication data across devices, highlighting the urgent need to reduce communication data volume in DLRM training.
However, as aforementioned, there are several challenges to addressing this issue due to the limitations of directly applying current error-bounded lossy compression techniques. To effectively tackle these hurdles, this paper focuses on the following key research questions: ❶ Error Bound Configuration: It’s essential to determine the optimal error bounds for various embedding tables across different training iterations to maintain training accuracy while enhancing compression ratios. ❷ DLRM-specific Lossy Compression Algorithm: There’s a critical need to devise a compression algorithm uniquely suited for DLRM’s embedding tables, aiming for elevated compression ratios without substantially compromising data integrity. ❸ GPU Compression Performance Optimization: Optimizing the GPU execution of our specialized compression and ensuring its smooth incorporation within DLRM training is crucial for improving overall training performance.
@@ III-A Overview of DLRM Training Pipeline with Compression
Offline Analysis. The purpose of this step is to obtain an optimized configuration by sampling and analyzing some iterations from the original training process. There are two tasks involved in offline analysis: Compressor Selection and Embedding Table Classification. In the Compressor Selection task, we evaluate various compressors using sampled data to select the best one for the current system. In the Embedding Table Classification task, we analyze the characteristics of embedding tables using sampled data and classify them according to different error-bound adjustment strategies.
@@ III-B Observation and In-Depth Analysis of Embedding Vector
TABLE I: Characteristics of their representative EMB tables from Criteo Kaggle dataset.
@@
Violently Vector
Homogenization
$\checkmark$ $\times$ $\times$
Gaussian
Distribution
$\checkmark$ $\checkmark$ $\times$
@@ III-C Adaptive Fine-Grain Error-Bound Adjust Strategy
Specifically, we divide the training period into two phases: an initial phase, characterized by rapid loss reduction, and a later phase, where the loss tends to stabilize. During the initial phase, we gradually decrease the error bound via a predefined decay function (e.g., logarithmic, stepwise) to minimize deviation from the original data, facilitating swift model convergence. In the later phase, we maintain a consistent error bound to ensure the model converges effectively. According to experiment results in Figure 5, it demonstrated that a step-wise (staircase descent) decay function offers the greatest compression benefits while ensuring model convergence. In that case, we select step-wise decay as the default decay function.
@@
Figure 6: EMB table sizes in Criteo Kaggle and Terabytes datasets.
@@
$\eta=\frac{N_{\text{original}}-N_{\text{compressed}}}{N_{\text{original}}}.$ (1)
@@
if c is 'large': EBConf[t] $\leftarrow$ LargeEB
if c is 'small': EBConf[t] $\leftarrow$ SmallEB
@@
func EMBClassification(hindex):
if hindex > S_EMB_hindex:
return 'small'
elif hindex < L_EMB_hindex:
return 'large'
else: return 'medium'
[…]
@@ III-D Optimized Compression Algorithm for DLRM Data
After identifying the optimal error bound for training, the next step involves fine-tuning the compressor within that error bound. Compressors typically aim to balance achieving a high compression ratio with maintaining high compression speed. The ideal compressor for our purposes should satisfy three criteria: ❶ Operate on the GPU to avoid data transfers between the device and host since embedding vectors reside on the GPU. ❷ Offer high compression throughput to minimize compression overhead and, consequently, accelerate overall DLRM training. ❸ Achieve a high compression ratio to maximize the benefits of reduced data volume during communication. To meet these requirements, we propose an optimized hybrid error-bounded lossy compression algorithm tailored specifically for DLRM data.
@@
$\text{speedup}=\frac{1}{\frac{1}{\text{CR}}+B\times(\frac{1}{T_{c}}+\frac{1}{T_{d}})}$ (2)
Overall, in our hybrid compression framework, we use this formula to pinpoint the most efficient compressor that maximizes speed-up for the given system configuration. We present the detail in Algorithm LABEL:alg:compressorSelection. In this pseudo code, Line 1 defines the compressor selection parameter. Line 2 defines the hyper-parameter for alternative compressors. The speedup calculation in line 6 refers to Equation (2). Although theoretically any compression algorithm could be included in our selection pool, for simplicity and effectiveness, we limit our final design to these two encoders.
@@
speedups $\leftarrow$ SpeedUpCompute(d, Compressors)
maxSup, bestCompressor = Max(speedups)
@@
[…]
Figure 7: Proposed buffer optimization resulting in single-kernel compression (top) and parallel decompression (bottom).
@@ III-E System Implementation and Performance Optimization
Compressor Fine-tuning. First, we refine our vector-based LZ compression algorithm, distinguishing it from the original LZ approach in two significant ways. ❶ Extended window size: Traditional LZ algorithms typically opt for a relatively small window size, like 4KB/8KB. However, given the longer repeated patterns observed in DLRM applications, with lengths ranging from 128-256 bytes (considering an embedding vector length of 32/64 with 32-bit floating point data type), we extend the window size to accommodate these patterns. ❷ Fixed pattern length: To minimize memory access and byte comparison time, we introduce fixed-length pattern matching. In contrast to the standard LZ process, which advances the pointer to the next byte in the absence of a match, our knowledge of the repeated pattern length allows us to leap forward many bytes in search of the next match, thereby avoiding fruitless comparisons and pointer movements. If the initial bytes of two embedding vectors differ, further comparison is unnecessary. This optimization also forces the compressor to match longer patterns rather than shorter patterns. Compared to the current state-of-the-art (SOTA) GPU lossless compressor nvCOMP-LZ4, our approach achieves 2.72$\times$ and 4.88$\times$ higher compression ratios on two datasets, respectively.
Buffer Optimization. Next, we implement multi-threading to minimize memory copy overhead and speed up decompression. Typically, compressors output compressed data to a memory chunk and return a pointer, a versatile but overhead-inducing solution. Given the all-to-all collective communication in DLRM training, where each data chunk must be compressed separately for transmission to each rank, this method introduces unnecessary memory copying since compressed data may not be stored contiguously. Besides, launching kernel multiple times introduces overhead of a lot of kernel launching. To address these, we optimize our compressor that not only introduces one kernel launching but also writes directly to the sending buffer. Figure 7 shows the workflow of compression and decompression in buffer optimization. In the compression process, we involve synchronization and Atomic Add to get the writing offset of each chunk. Furthermore, we leverage multi-threading for compression and decompression, partitioning the compressed data into multiple chunks for simultaneous processing. Although GPU compressors typically achieve high resource utilization, some operations cannot fully leverage GPU resources. Thus, executing multiple decompression kernels in parallel is faster than serial decompression.
@@ IV-A Experimental Setup
TAB. ID EB.
# Ori.
Patterns
# Quant.
Patterns
Batch Size Homo Index
@@
TAB. ID EB.
# Ori.
Patterns
# Quant.
Patterns
Batch Size Homo Index
@@ IV-B Evaluation of Error Bound Adjustment Strategy
To assess the effects of lossy compression on model accuracy, we evaluate the model prediction accuracy of our training method compared with original DLRM training using both FP32 and FP16 precisions, alongside a SOTA low precision approach employing 8-bit quantization. Current standards deem an accuracy loss within 0.02% as acceptable in production models [0.02threshold].
@@
Figure 8(a) and figure 8(b) display a comparison of accuracy convergence curves and accuracy delta curves across these methods, respectively. Through extensive experimentation, we determine and apply a suitable fixed global error bound (i.e., 0.02) for each model to guarantee convergence. On the two datasets we utilized, the average prediction accuracy losses are 0.0031% and 0.0042%, respectively.
@@
As illustrated in Table II, embedding tables are classified into three categories according to their Homogenization Index scores, with assigned error bounds of 0.01, 0.03, and 0.05, respectively. The accuracy evaluation, detailed in Figure 9(a), reveals that our approach, which applies specific error bounds to different tables rather than a uniform global error bound, maintains the model’s accuracy intact. Additionally, this method achieves a higher compression ratio, up to 1.21$\times$ on the Criteo Kaggle dataset, compared to the fixed global error bound strategy.
@@
Figure 9: Accuracy and compression ratio of our method with proposed table-wise EB configuration strategy on different datasets. e.g. subgraph(b) represents compression ratio on embedding table 0
(a) Accuracy
(b) Compression Ratio
Figure 10: Accuracy and compression ratio of our method with two decay methods on different datasets. Decay__\__2x represents the error bound decayed from 2 times base error bound to base error bound; Drop__\__2x represents the error bound dropped from 2 times base error to base error bound at the end of the initial phase. E.g. subgraph (b) represents the compression ratio on embedding table 1.
@@
As we discussed in III-C. This experiment’s results of different compressors led us to employ the step-wise function for comparison against the aggressive adjustment approach. Comparative experiments illustrated in Figure 10 reveal how the model training and compression ratios are affected when the error bound is reduced from twice and three times the conservative error bound down to the conservative error bound.